OpenStudy (anonymous):

What is the exact value of tan (arcsin ( 2/5))

OpenStudy (anonymous):

\[\frac{ 2 }{ \sqrt{21} }\]

OpenStudy (anonymous):

thanks, how'd you get that answer?

OpenStudy (anonymous):



OpenStudy (anonymous):

The tangent is:\[\frac{2}{\sqrt{21}} \]

OpenStudy (anonymous):

Thank you